# Semi-empirical Modeling of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

### v1

@@ 2.2 Modeling the growth by mergers
where $K(M_{\bullet},M_{\bullet}^{\prime},t)$ is a time-dependent coagulation kernel, which has dimension of inverse of a time (i.e., a merging rate); the kernel is routinely taken to be symmetric in its mass dependencies, i.e. $K(M_{\bullet},M_{\bullet}^{\prime},t)=K(M_{\bullet}^{\prime},M_{\bullet},t)$. The first integral on the right hand side of the above equation is a production term which refers to a merging event between masses $M_{\bullet}^{\prime}$ and $M_{\bullet}-M_{\bullet}^{\prime}$ to produce the mass $M_{\bullet}$. The factor $1/2$ avoids over-counting since the labels of the two merging masses may be interchanged given the symmetric nature of the kernel; it could be removed by changing the upper limit of integration to $M_{\bullet}/2$. The second addendum is a destruction term which accounts for merging between the mass $M_{\bullet}$ and a companion with mass $M_{\bullet}^{\prime}$. We stress that the Smoluchowski equation is only appropriate when the population evolution is dominated by binary aggregations, and instead should not be employed when the probability of multiple interactions is appreciable. For example, it is not very apt to describe the merging history of dark matter halos at high redshift (or during any fast accretion phase), since the kernels inferred from the excursion set formalism or extracted from numerical simulations imply that the average number of halo mergers is often larger than two (e.g., [144, 147]). In the present context, we make the reasonable hypothesis that the great majority of the supermassive BH mergers are in fact binary ones and that multiple interactions are possible but highly improbable.
@@
where $p$ is an integer (e.g., [141, 142, 143]). The above equation makes evident that $\partial_{t}\langle M_{\bullet}\rangle=0$ and $\partial_{t}\langle M_{\bullet}^{2}\rangle>0$; in other words, the coagulation processes does not cause a drift in the mass distribution, but just enhances its dispersion by redistributing masses from the low to the high mass end. We mention that there is an exception to the mass conservation $\langle\dot{M}_{\bullet}\rangle=0$ found above: when the kernel dependence on the mass is sufficiently strong, like in the well-known case of a multiplicative kernel $K(M_{\bullet},M_{\bullet}^{\prime})\propto M_{\bullet}\,M_{\bullet}^{\prime}$, then a sort of phase transition occurs, in that some mass may be lost in a finite time due to the appearance of clusters with infinite mass [148]. Such a phenomenon is called ‘gelation’ and has attracted a lot of interest in the scientific community for its practical applications to the physics of aereosol, polymers and molecular dynamics; however, this is not of our interest here (for a review, see [141]).
@@
Note that the quantity $K(M_{\bullet,1},M_{\bullet,2})\,N(M_{\bullet,1})\,N(M_{\bullet,2})\equiv{\rm d}^{2}\mathcal{R}/{\rm d}M_{\bullet,1}\,{\rm d}M_{\bullet,2}$ just represents a merging rate density per bins of the two merging masses $M_{\bullet,1}$ and $M_{\bullet,2}$. However, in the present context it is more convenient to work with the merger rate density ${\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}/{\rm d}M_{\bullet,1+2}\,{\rm d}q$ per bin of total (descendant) mass $M_{\bullet,1+2}\equiv M_{\bullet,1}+M_{\bullet,2}$ and mass ratio $q\equiv\min(M_{\rm\bullet,1};M_{\bullet,2})$ $/\max(M_{\rm\bullet,1};M_{\bullet,2})$, i.e. the ratio between the lesser and the more massive merging BHs (by definition). This can be achieved by a suitable change of variables in each term of the Smoluchowski equation, taking into account that the the Jacobian of the transformation is just $M_{\bullet,1+2}/(1+q)^{2}$. Identifying the two appropriate merging masses in each integral of Equation (2.8), and splitting the destruction term into two parts to always have mass ratios $q<1$, the overall Smoluchowski equation (2.8) can be rewritten in a rather compact form as (see also [108])
$[\partial_{t}N(M_{\bullet},t)]_{\bullet\bullet\rightarrow\bullet}=\int_{0}^{1}{\rm d}q\,\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big{|}M_{\bullet}}-\int_{0}^{1}{\rm d}q\,(1+q)\,\left[\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big{|}M_{\bullet}\,(1+q)}+q\,\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big{|}M_{\bullet}\,(1+q)/q}\right]\;.$ (2.10)
@@
Simulations indicate that the dependencies on mass, mass ratio and redshift of the merger rate densities per descendant dark matter halo or per descendant galaxy can be factorized [149, 150], so we make the simplifying assumption that the same holds for the hosted supermassive BHs. Therefore we parameterize the merger rate density per BH as
@@
∙∙→∙ $=\mathcal{R}_{\bullet\bullet\rightarrow\bullet}(t)\,\left(\frac{M_{\bullet}}{10^{10}\,M_{\odot}}\right)^{\eta}\,\Bigg{\{}N(M_{\bullet},t)+$ (2.12)
$-\int_{0}^{1}{\rm d}q\;p(q)\,(1+q)^{1+\eta}\,\left[N(M_{\bullet},t)_{\big{|}M_{\bullet}\,(1+q)}\,+q^{1-\eta}\,N(M_{\bullet},t)_{\big{|}M_{\bullet}\,(1+q)/q}\right]\Bigg{\}}\;.$
We parameterize the redshift-dependent merger rate normalization in terms of a rescaling of the galaxy merger rate as estimated from galaxy pair counts observed by JWST (see [151, 152, 153]). Specifically, we take
@@
in terms of a scaling factor $f_{\bullet\bullet\rightarrow\bullet}$ and of the parameters $\mathcal{R}_{0}\approx 0.013$ Gyr-1, $\omega\approx 3.4$ and $\tau\approx 0.14$ set by [152]. This choice is appropriate since it would be impossible to constrain the detailed redshift dependence of the BH merger rate relying only on the data concerning BH demographics and clustering, and on the mild integrated redshift dependence of the stochastic background in GWs (see below).
@@
In addition, note that in the above we are neglecting any possible time delay between the merger of the BHs and that of the host galaxies, which constitutes the standard assumption in the analysis of the GW background [103, 154]; to wit, this maximizes the contribution of binary mergers. Finally, following [110, 107] we take the mass ratio distribution $p(q)\propto q^{-\delta}\,\theta_{\rm H}[q\geq q_{\rm min}]$ as a power-law with index $\delta$ extending down to a minimum mass ratio $q_{\rm min}$, held fixed to the reference values $\delta=-1$ and $q_{\rm min}=0.1$. In fact, the dependence of the results on the mass ratio distribution is very weak, so there is no point in trying to fit the related parameters.
2.3 Modeling BH demographics
The BH demographics, i.e. the BH mass function at any cosmic time, can be reconstructed from the continuity (plus Smoluchowsi) equation expressed as:
@@ 2.3 Modeling BH demographics
Fiducially, we initialize the computation at $z_{\rm ini}\sim 7$ which is the maximal redshift at which measurements of the AGN luminosity function by [74] do not require extrapolation, and there we assume as initial condition a null active BH mass function. However, we also try to extend the initial redshift out to $z_{\rm ini}\sim 10$ by extrapolating the observed AGN luminosity functions. We also try to vary the initial condition by assuming a duty cycle equal to one and a narrowly peaked mass function or the mass function stemming from various light/seed population models [155, 156]. Basically, we have found that as for the determination of the BH mass function at $z\lesssim 6$ and below, the initial redshift and any reasonable initial conditions are marginally relevant since the accumulated mass from accretion soon largely exceeds that in the original BH population, and the impact of BH mergers on the latter at high redshift will turn out to be relatively minor.
2.4 Modeling BH Clustering
A further observable to better constrain the parameters of our semi-empirical model, and in particular to contribute in reducing the degeneracy between the radiative efficiency and Eddington ratios is to exploit the large-scale BH clustering [23, 56]. Specifically, we will be interested in comparing our model outcomes to data concerning the bias $b(M_{\bullet},z)$, as provided at relatively low redshifts $z\lesssim 0.25$ by optical and X-ray observations of AGNs with robust BH mass determinations [157, 158].
From the modeling point of view, to compute the BH clustering, we start from the large-scale bias $b(M_{\rm H},z)$ of dark matter (DM) halos with mass $M_{\rm H}$, which can be easily extracted from $N-$body simulations (see [159]). To translate the halo into BH masses, we exploit the standard abundance matching technique [122, 160], in the version by [24] that includes the effect of the scatter in the halo-to-BH mass relation. This procedure is also supported by the recent pairwise residuals analysis by [161], which has highlighted that the correlation between BH and host halo mass appears as strong as the one with velocity dispersion, as corroborated by direct fitting (e.g., [162]) and clustering analysis (e.g., [163]).
Specifically, we obtain $M_{\rm H}(M_{\bullet},z)$ by requiring the cumulative number densities in BHs and halos to match as:
@@ 2.4 Modeling BH Clustering
where the scatter of the relation $\sigma_{\log M_{\rm H}}$ dex is a fitting parameter. The integrand on the left hand side encases the BH mass function, that we obtain by solving the continuity equation. The right hand side involves the CDM halo mass function $N(\log M_{\rm H},z)$, that we take from [164]. In fact, we compute numerically both the halo mass function [^6] We also include the correction for sub-halos in the halo mass function. However, this becomes relevant only at halo masses $M_{\rm H}\lesssim 10^{12}\,M_{\odot}$ and thus it is found to play a minor role in the present context. and the halo bias via the COLOSSUS package [165]. Finally, the halo bias as a function of BH mass $b(M_{\bullet},z)$, needed to compare with data, is then obtained by convolving the bias as a function of halo mass with the halo mass vs. BH mass relation, properly taking into account its dispersion $\sigma_{\log M_{\rm H}}$.
2.5 Modeling the stochastic GW background
Binary mergers of supermassive BH throughout cosmic history can produce a stochastic backround of GW with typical nano-Hz frequency, that is within the reach of present PTA experiments. The characteristic strain $h_{c}$ of the stochastic GW background from binary BH mergers (in circular orbits) as a function of the observed frequency $f$ is given by [166]
@@ 2.5 Modeling the stochastic GW background
Note that very often results of nano-Hz GW experiments are quoted in terms of a normalization $\mathcal{A}$ for the characteristic strain [167] written in the form $h_{c}(f)=\mathcal{A}\,(f/{\rm yr}^{-1})^{-2/3}$, where the constant $\mathcal{A}$ can be computed easily from the expression above. The $f^{-2/3}$ scaling holds for circular orbits, and it can actually be altered by eccentricity or by environmental effects (in that they tend to somewhat reduce the strain and flatten its frequency dependence; see [168]), but these go beyond the scope of the present paper.
@@ 3 Data and analysis
We now turn to constrain the parameters of our semi-empirical model by comparing its outcomes with several robust datasets. Specifically, we exploit local determinations of the BH mass function, of the BH clustering, and a recent estimate of the nano-Hz GW background.
@@
As mentioned in Section 1, the local BH mass function is routinely obtained via the convolution of a galaxy statistics with a scaling relationship between BH mass and galaxy properties. We choose to consider three cases. One is obtained by combining the local stellar mass function or SMF $N(M_{\star})$ by [169] with the $M_{\bullet}-M_{\star}$ relationship for early-type galaxies by [170, 56]. Another one is obtained by combining the local galaxy velocity dispersion function or VDF $N(\sigma)$ by [171] and the observed $M_{\bullet}-\sigma$ relationship by [48]. Finally, the last one is obtained by combining the VDF with the ‘debiased’ version of the $M_{\bullet}-\sigma$ by [54]; this is meant to constitute a solid lower limit to the local demography, as pointed out by the recent results in [161]. Such binned local BH mass functions are reported in Table 1. Our analysis will be conducted independently for each of these three cases, which confidently bracket the overall systematic uncertainty [^7] Note that using the SMF and the intrinsic $M_{\bullet}-M_{\star}$ relationship by [54] produces a local BH mass function consistent with the combination VDF + intrinsic $M_{\bullet}-\sigma$, as shown explicitly by [56].. They also span the range covered by other literature estimates from different galaxy statistics and/or relationships; for example, the recent BH mass function by [172] strikes an intermediate course among the cases based on the observed $M_{\bullet}-\sigma$ and $M_{\bullet}-M_{\star}$ relations (see also Figure 2). As to BH clustering, we rely on the measurements of the large-scale bias $b(M_{\bullet})$ as a function of BH mass at $z\sim 0.25$ by [157, 158] and at $z\sim 0.04$ by [173, 174]. As to the GW background, we base on the most recent estimate from PTA by the NANOGrav collaboration [96], in terms of the strain amplitude parameter $\mathcal{A}=(2.4\pm 0.7)\times 10^{-15}$ expressed by Equation (2.16). We stress again that an added value of the present work is in attempting to fit jointly all these observables. In Appendix B we will also test our model outcomes when excluding the PTA measurement from the likelihood, i.e. fitting only for BH demographics and clustering and then compare the resulting GW background with data.
@@
We will also confront the outcomes of our semi-empirical model with other auxiliary datasets not included in the fitting procedure: measurements of the local duty cycle $\langle\delta\rangle(M_{\bullet})$ of AGN activity as a function of BH mass (taking into account appropriate luminosity threshold) by [175, 176, 138]; estimates of the active BH mass function at high-redshift $z\sim 4-6$, including BAL quasars and little red dots by [68, 90, 89, 88, 177, 91]; BH vs. halo mass relationship inferred from different methods [178, 157, 179, 180]; evolution with redshift of the average Eddington ratio for various kinds of AGNs [65, 181, 182, 37, 89, 90, 39, 84, 183, 184, 88]. Being rather uncertain and subject to large systematics or observational biases, these datasets are not employed to fit the parameters of our semi-empirical model, but rather are exploited as a-posteriori validation.
$\log M_{\bullet}$ [$M_{\odot}$] $\log N(\log M_{\bullet})$ [Mpc-1 dex-1]
@@
$7.00$ $-2.71\pm 0.06$ $-2.59\pm 0.04$ $-2.70\pm 0.05$
$7.50$ $-2.60\pm 0.06$ $-2.69\pm 0.04$ $-2.88\pm 0.05$
$8.00$ $-2.62\pm 0.08$ $-2.87\pm 0.05$ $-3.18\pm 0.05$
$8.50$ $-2.79\pm 0.10$ $-3.17\pm 0.06$ $-3.65\pm 0.07$
$9.00$ $-3.18\pm 0.12$ $-3.63\pm 0.07$ $-4.34\pm 0.11$
$9.50$ $-3.83\pm 0.16$ $-4.32\pm 0.11$ $-5.30\pm 0.18$
$10.0$ $-4.76\pm 0.23$ $-5.28\pm 0.18$ $-6.59\pm 0.28$
Table 1: The local BH mass functions exploited in our analysis. These include the determination from the stellar mass function (SMF) from [169] plus the $M_{\bullet}-M_{\star}$ relationship from [170, 56] , the velocity dispersion function (VDF) from [171] plus the $M_{\bullet}-\sigma$ relationship in its observed [obs.] version from [48] or in its debiased [deb.] version from [54]. See text for more details.
@@
We sample the parameter posterior distributions $\mathcal{P}(\theta)\propto\mathcal{L}(\theta)\,\pi(\theta)$ via the MCMC Python package emcee [185], running it with $10^{5}$ steps and $100$ walkers; each walker is initialized with a random position extracted from the priors discussed above. To speed up the computation, the active BH mass function is pre-determined solving Equation (2.5) on a grid of redshift and Eddington ratio, and then its parameters are interpolated on the fly during the MCMC run. To more efficiently sample the parameter space, we adopt a mixture of differential evolution and snooker moves of the walkers, in proportion of $0.8$ and $0.2$ respectively, that emulates a parallel tempering algorithm. After checking the auto-correlation time, we remove the first $50\%$ of the flattened chain to ensure burn-in.

### v2

@@ 2.3 Growth by mergers
where $K(M_{\bullet},M_{\bullet}^{\prime},t)$ is a time-dependent coagulation kernel, which has dimension of inverse of a time (i.e., a merging rate); the kernel is routinely taken to be symmetric in its mass dependencies, i.e. $K(M_{\bullet},M_{\bullet}^{\prime},t)=K(M_{\bullet}^{\prime},M_{\bullet},t)$. The first integral on the right hand side of the above equation is a production term which refers to a merging event between masses $M_{\bullet}^{\prime}$ and $M_{\bullet}-M_{\bullet}^{\prime}$ to produce the mass $M_{\bullet}$. The factor $1/2$ avoids over-counting since the labels of the two merging masses may be interchanged given the symmetric nature of the kernel; it could be removed by changing the upper limit of integration to $M_{\bullet}/2$. The second addendum is a destruction term which accounts for merging between the mass $M_{\bullet}$ and a companion with mass $M_{\bullet}^{\prime}$.
We stress that the Smoluchowski equation is only appropriate when the population evolution is dominated by binary aggregations, and instead should not be employed when the probability of multiple interactions is appreciable. For example, it is not very apt to describe the merging history of dark matter halos at high redshift (or during any fast accretion phase), since the kernels inferred from the excursion set formalism or extracted from numerical simulations imply that the average number of halo mergers is often larger than two (e.g., [Benson2005, Neistein2008]). However, for supermassive BHs the binary merger assumption is more reasonable. In fact, the issue has been somewhat investigated via aimed dynamical simulations [AmaroSeoane2023, Koehn2023, Sayeb2024]. These show that only a minority around $10-20\%$ of binary supermassive BHs can form a triple system by capturing another BH intruder. Moreover, only a fraction $\lesssim 5\%$ actually form a triple systems at parsec scale experiencing chaotic three-body dynamical interaction; in the remaining majority of the cases two of the three BHs merge first and the third is ejected or left orbiting at large separation, so that the the merger is still binary. Thus despite merging triple BH systems could be interesting for future GW detector like LISA since they are expected to feature some peculiar signatures, their overall number is very small, and their contribution to the stochastic GW background irrelevant, compared to the standard binary merger channel.
@@
where $p$ is an integer (e.g., [Leyvraz2003, Smoluchowski1916, Cavaliere1992]). The above equation makes evident that $\partial_{t}\langle M_{\bullet}\rangle=0$ and $\partial_{t}\langle M_{\bullet}^{2}\rangle>0$; in other words, the coagulation processes does not cause a drift in the mass distribution, but just enhances its dispersion by redistributing masses from the low to the high mass end. We mention that there is an exception to the mass conservation $\langle\dot{M}_{\bullet}\rangle=0$ found above: when the kernel dependence on the mass is sufficiently strong, like in the well-known case of a multiplicative kernel $K(M_{\bullet},M_{\bullet}^{\prime})\propto M_{\bullet}\,M_{\bullet}^{\prime}$, then a sort of phase transition occurs, in that some mass may be lost in a finite time due to the appearance of clusters with infinite mass [Cavaliere1991]. Such a phenomenon is called ‘gelation’ and has attracted a lot of interest in the scientific community for its practical applications to the physics of aereosol, polymers and molecular dynamics; however, this is not of our concern here (for a review, see [Leyvraz2003]).
@@
Note that the quantity $K(M_{\bullet,1},M_{\bullet,2})\,N(M_{\bullet,1})\,N(M_{\bullet,2})\equiv{\rm d}^{2}\mathcal{R}/{\rm d}M_{\bullet,1}\,{\rm d}M_{\bullet,2}$ just represents a merging rate density per bins of the two merging masses $M_{\bullet,1}$ and $M_{\bullet,2}$. However, in the present context it is more convenient to work with the merger rate density ${\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}/{\rm d}M_{\bullet,1+2}\,{\rm d}q$ per bin of total (descendant) mass $M_{\bullet,1+2}\equiv M_{\bullet,1}+M_{\bullet,2}$ and mass ratio $q\equiv\min(M_{\rm\bullet,1};M_{\bullet,2})$ $\max(M_{\rm\bullet,1};M_{\bullet,2})$, i.e. the ratio between the lesser and the more massive merging BHs (by definition). This can be achieved by a suitable change of variables in each term of the Smoluchowski equation, taking into account that the the Jacobian of the transformation is just $M_{\bullet,1+2}/(1+q)^{2}$. Identifying the two appropriate merging masses in each integral of Equation (2.10), and splitting the destruction term into two parts to always have mass ratios $q<1$, the overall Smoluchowski equation (2.10) can be rewritten in a rather compact form as (see also [SatoPolito2025])
${}_{\bullet\bullet\rightarrow\bullet}=\int_{0}^{1}{\rm d}q\,\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big|M_{\bullet}}$ $-\int_{0}^{1}{\rm d}q\,(1+q)\,\left[\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big|M_{\bullet}\,(1+q)}+\right$. (2.12)
$\left.+q\,\frac{{\rm d}^{2}\mathcal{R}_{\bullet\bullet\rightarrow\bullet}}{{\rm d}M_{\bullet}\,{\rm d}q}_{\big|M_{\bullet}\,(1+q)/q}\right]\;.$
@@
Simulations indicate that the dependencies on mass, mass ratio and redshift of the merger rate densities per descendant dark matter halo or per descendant galaxy can be factorized [Fakhouri2009, Rodriguez2016], so we make the simplifying assumption that the same holds for the hosted supermassive BHs. Therefore we parameterize the merger rate density per BH as
@@
${}_{\bullet\bullet\rightarrow\bullet}=\mathcal{R}_{\bullet\bullet\rightarrow\bullet}(t)\,\left(\frac{M_{\bullet}}{10^{10}\,M_{\odot}}\right)^{\eta}$ $\Bigg\{N(M_{\bullet},t)-\int_{0}^{1}{\rm d}q\;p(q)\,(1+q)^{1+\eta}\,\left[N(M_{\bullet},t)_{\big|M_{\bullet}\,(1+q)}\right$. (2.14)
$\left.+q^{1-\eta}\,N(M_{\bullet},t)_{\big|M_{\bullet}\,(1+q)/q}\right]\Bigg\}\;.$
We parameterize the redshift-dependent merger rate normalization in terms of a rescaling of the galaxy merger rate as estimated from galaxy pair counts observed by JWST (see [Casteels2014, Duan2025, Puskas2025]). Specifically, we take
@@
in terms of a scaling factor $f_{\bullet\bullet\rightarrow\bullet}$ and of the parameters $\mathcal{R}_{0}\approx 0.013$ Gyr-1, $\omega\approx 3.4$ and $\tau\approx 0.14$ set by [Duan2025]. This choice is appropriate since it would be impossible to constrain the detailed redshift dependence of the BH merger rate relying only on the data concerning BH demographics and clustering, and on the mild integrated redshift dependence of the stochastic background in GWs (see below).
@@
In addition, note that in the above we are neglecting any possible time delay between the merger of the BHs and that of the host galaxies, which constitutes the standard assumption in the analysis of the GW background [Sesana2016, Agazie2023b]; to wit, this maximizes the contribution of binary mergers.
Finally, following [Kelley2017, SatoPolito2024] we take the mass ratio distribution $p(q)\propto q^{-\delta}\,\theta_{\rm H}[q\geq q_{\rm min}]$ as a power-law with index $\delta$ extending down to a minimum mass ratio $q_{\rm min}$, held fixed to the reference values $\delta=-1$ and $q_{\rm min}=0.1$. In fact, the dependence of the results on the mass ratio distribution is very weak, so there is no point in trying to fit the related parameters.
2.4 BH demographics
The direct output of the continuity equation is constituted by the BH demographics, i.e. the BH mass function at any cosmic time. This is found by solving the continuity equation:
@@ 2.4 BH demographics
Fiducially, we initialize the computation at $z_{\rm ini}\sim 7$ which is the maximal redshift at which measurements of the AGN luminosity function by [Shen2020] do not require extrapolation, and there we assume as initial condition a mean duty cycle $\langle\delta\rangle(M_{\bullet},z_{\rm ini})=0.5$ independent of mass [Shankar2013]. Then by definition of the duty cycle one has $N_{\rm AGN}(\log M_{\bullet},z_{\rm ini})=N(\log M_{\bullet},z_{\rm ini})\,\langle\delta\rangle=N(\log M_{\bullet},z_{\rm ini})/2$ and hence, given the observed AGN luminosity function $N(L_{\rm AGN},z_{\rm ini})$, one can determine the initial BH mass function $N(\log M_{\bullet},z_{\rm ini})$ just by solving Equation (2.6). To explore the dependence of our results on the initial condition, we try to use $z_{\rm ini}\sim 10$ by extrapolating the observed AGN luminosity functions. We also try to assume a different initial duty cycle in the range $\langle\delta\rangle(M_{\bullet},z_{\rm ini})=0.1-1$. Finally, we try a narrowly peaked mass function or the mass function stemming from various light-seed population models [Volonteri2010, Sicilia2022a]. Basically, we have found that as for the determination of the BH mass function at $z\lesssim 4$ and below, the initial redshift and reasonable initial conditions are marginally relevant since the accumulated mass from accretion soon largely exceeds that in the original BH population (see also [Aversa2015, Sicilia2022, Roberts2025]), and the impact of BH mergers on the latter at high redshift will turn out to be relatively minor.
2.5 BH Clustering
A further observable to better constrain the parameters of our semi-empirical framework, and in particular to contribute in reducing the degeneracy between the parameters is to exploit the large-scale BH clustering [Shankar2013, Shankar2020]. Specifically, we will be interested in comparing our model outcomes to data concerning the bias $b(M_{\bullet},z)$, as provided at relatively low redshifts $z\lesssim 0.25$ by optical and X-ray observations of AGNs with robust BH mass determinations [Krumpe2015, Krumpe2023].
In the context of our semi-empirical framework, the BH clustering signal can be computed by combining the BH mass function output from the continuity equation to the mass function and bias of dark matter halos as extracted from NN-body simulations. In more detail, the procedure is the following. First of all, we employ the standard abundance matching technique [Shankar2004, Moster2013] to relate BH and halo masses in a nearly model-independent way, the only assumption being the existence of a monotonic mapping $M_{\rm H}(M_{\bullet},z)$ between the two variables, at least on statistical grounds. In particular, we use the version of abundance matching by [Aversa2015] that includes the effect of the scatter in the halo-to-BH mass relation. This boils down to require the cumulative number densities in BHs and halos to match as:
@@ 2.5 BH Clustering
[…]
2.6 Stochastic GW background
Binary mergers of supermassive BH throughout cosmic history can produce a stochastic backround of GW with typical nano-Hz frequency, that is within the reach of present PTA experiments. In the context of our semi-empirical framework, this can be easily computed from the outputs of the continuity equation, and in particular from the BH mass function and the BH merger rate. Specifically, the characteristic strain $h_{c}$ of the stochastic GW background from binary BH mergers (in circular orbits) as a function of the observed frequency $f$ is given by [Phinney2001]
@@ 2.6 Stochastic GW background
where $N(M_{\bullet},t)$ is the BH mass function obtained from solving the continuity Equation (2.16) and $\mathcal{R}_{\bullet\bullet\rightarrow\bullet}$ is given by Equation (2.15). Note that very often results of nano-Hz GW experiments are quoted in terms of a normalization $\mathcal{A}$ for the characteristic strain [Jenet2006] written in the form $h_{c}(f)=\mathcal{A}\,(f/{\rm yr}^{-1})^{-2/3}$, where the constant $\mathcal{A}$ can be computed easily from the expression above. The $f^{-2/3}$ scaling holds for circular orbits, and it can actually be altered by eccentricity or by environmental effects (in that they tend to somewhat reduce the strain and flatten its frequency dependence; see [Fastidio2025]), but these go beyond the scope of the present paper.
@@ 3 Data and analysis
We now turn to constrain the parameters of our semi-empirical framework by comparing its outcomes with various datasets. Specifically, we exploit local determinations of the BH mass function, of the BH clustering, and a recent estimate of the nano-Hz GW background.
@@
As mentioned in Section 1, the local BH mass function cannot be estimated directly, since the number of robustly measured BH masses is small (less than a hundred). Thus this is routinely obtained via the convolution of a galaxy statistics (measured on a much large sample of galaxies) with a scaling relationship between BH mass and galaxy properties (calibrated on the small sample of directly measured BH masses). Unfortunately, this procedure brings in a certain degree of uncertainty in the determination of the local BH mass function, which is important to take into account. To this purpose, we consider three cases. In the first case, the local BH mass function is obtained by combining the local stellar mass function or SMF $N(M_{\star})$ by [Moffett2016] with the $M_{\bullet}-M_{\star}$ relationship for early-type galaxies by [Savorgnan2016, Shankar2020]. In the second case, the local BH mass function is obtained by combining the local galaxy velocity dispersion function or VDF $N(\sigma)$ by [Sheth2003] and the observed $M_{\bullet}-\sigma$ relationship by [McConnell2013]. Finally, in the third case the local BH mass function is obtained by combining the VDF with the ‘debiased’ version of the $M_{\bullet}-\sigma$ by [Shankar2016]; this is meant to constitute a solid lower limit to the local demography, as pointed out by the recent results in [Shankar2025]. Such binned local BH mass functions are reported in Table 1. Our analysis will be conducted independently for each of these three cases, which confidently bracket the overall systematic uncertainty [^8] Note that using the SMF and the intrinsic $M_{\bullet}-M_{\star}$ relationship by [Shankar2016] produces a local BH mass function consistent with the combination VDF + intrinsic $M_{\bullet}-\sigma$, as shown explicitly by [Shankar2020].. They also span the range covered by other literature estimates from different galaxy statistics and/or relationships. For example, the recent BH mass function by [Liepold2024] strikes an intermediate course between our cases based on the observed $M_{\bullet}-\sigma$ and $M_{\bullet}-M_{\star}$ relations (see dotted line in Figure 4). As another example, the determination by [Shen2020], obtained convolving the AGN luminosity function with the AGN luminosity vs. Eddington ratio relationship from SDSS [Shen2009, Nobuta2012] and further correcting for obscured AGNs and inactive BHs, is in between our cases based on the observed and the debiased $M_{\bullet}-\sigma$ relations (see shaded area in Figure 4).
As to BH clustering, we rely on the measurements of the large-scale bias $b(M_{\bullet})$ as a function of BH mass at $z\sim 0.25$ by [Krumpe2015, Krumpe2023] and at $z\sim 0.04$ by [Powell2018, Allevato2021]. Note that a wealth of additional datasets (e.g., [Shen2009, Allevato2011, White2012, Shen2013, Allevato2014, Allevato2016, Mountrichas2016, Magliocchetti2017, Hale2018, Powell2018, Eftekharzadeh2019, Ross2020, Krishnan2020, Mazumder2022, Hale2025]) are available at different redshifts for the clustering of AGNs as a function of luminosity from optical, X-ray and radio selected sample. However, we cannot rely on these datasets to independently test our framework. This is because we used as an input the AGN luminosity function, hence the relation between AGN luminosity and halo mass (obtained via abundance matching, see Section 2.5) and the ensuing clustering signal as a function of AGN luminosity is reproduced by construction. Focusing instead on the clustering signal as a function of BH mass can independently constrain the halo to BH mass relationship and hence ultimately the BH mass function, which is the main output of our framework (see also [Shankar2020, Allevato2021]).
As to the GW background, we rely on the most recent estimate from PTA by the NANOGrav collaboration [Agazie2023], in terms of the strain amplitude parameter $\mathcal{A}=(2.4\pm 0.7)\times 10^{-15}$ expressed by Equation (2.18). We stress again that an added value of the present work is in attempting to fit jointly all these observables. In Appendix B we will also test our model outcomes when excluding the PTA measurement from the likelihood, i.e. fitting only for BH demographics and clustering and then compare the resulting GW background with data.
@@
We will also confront the outcomes of our semi-empirical framework with other auxiliary datasets not included in the fitting procedure: measurements of the local duty cycle $\langle\delta\rangle(M_{\bullet})$ of AGN activity as a function of BH mass (taking into account appropriate luminosity threshold) by [Kauffmann2003, Goulding2010, Schulze2010]; estimates of the active BH mass function at high-redshift $z\sim 4-6$, including BAL quasars and little red dots by [Wu2022, Lai2024, He2024, Matthee2024, Taylor2025, Kokorev2024, Geris2025]; BH vs. halo mass relationship inferred from different methods [Marasco2021, Krumpe2015, Robinson2021, Li2024]; evolution with redshift of the average Eddington ratio for various kinds of AGNs [Vestergaard2009, Nobuta2012, Dai2014, Farina2022, He2024, Lai2024, Fan2023, Harikane2023, Greene2024, Maiolino2024, Matthee2024]. Being affected by appreciable uncertainties and possibly by systematics or observational biases, these datasets are not employed to fit the parameters of our semi-empirical framework, but rather are exploited as a-posteriori validation.
$\log M_{\bullet}$ $M_{\odot}$] $\log N(\log M_{\bullet})$ [Mpc-1 dex-1]
@@
$7.00$ $2.71\pm 0.06$ $2.59\pm 0.04$ $2.70\pm 0.05$
$7.50$ $2.60\pm 0.06$ $2.69\pm 0.04$ $2.88\pm 0.05$
$8.00$ $2.62\pm 0.08$ $2.87\pm 0.05$ $3.18\pm 0.05$
$8.50$ $2.79\pm 0.10$ $3.17\pm 0.06$ $3.65\pm 0.07$
$9.00$ $3.18\pm 0.12$ $3.63\pm 0.07$ $4.34\pm 0.11$
$9.50$ $3.83\pm 0.16$ $4.32\pm 0.11$ $5.30\pm 0.18$
$10.0$ $4.76\pm 0.23$ $5.28\pm 0.18$ $6.59\pm 0.28$
Table 1: The local BH mass functions exploited in our analysis. These include the determination from the stellar mass function (SMF) from [Moffett2016] plus the $M_{\bullet}-M_{\star}$ relationship from [Savorgnan2016, Shankar2020] , the velocity dispersion function (VDF) from [Sheth2003] plus the $M_{\bullet}-\sigma$ relationship in its observed [obs.] version from [McConnell2013] or in its debiased [deb.] version from [Shankar2016]. See text for more details.
@@
We sample the parameter posterior distributions $\mathcal{P}(\theta)\propto\mathcal{L}(\theta)\,\pi(\theta)$ via the MCMC Python package emcee [emcee], running it with $10^{5}$ steps and $100$ walkers; each walker is initialized with a random position extracted from the priors discussed above. To speed up the computation, the active BH mass function is pre-determined solving Equation (2.8) on a grid of redshift and Eddington ratio, and then its parameters are interpolated on the fly during the MCMC run. To more efficiently sample the parameter space, we adopt a mixture of differential evolution and snooker moves of the walkers, in proportion of $0.8$ and $0.2$ respectively, that emulates a parallel tempering algorithm. After checking the auto-correlation time, we remove the first $50\%$ of the flattened chain to ensure burn-in.